OpenStudy (anonymous):

Write the expression as the sine, cosine, or tangent of an angle. sin 9x cos x - cos 9x sin x

OpenStudy (anonymous):

sin(8x)

OpenStudy (anonymous):

Thank you @asaad123

OpenStudy (anonymous):

How did you get the answer?

OpenStudy (anonymous):

since sin u cos v -cos u sin v =sin (u-v) sin(9x-x)=sin(8x)